OpenStudy (anonymous):

Write the quotient in the form a+bi 1+3i over 5+3i

OpenStudy (anonymous):

\[\frac{1+3i}{5+3i}=\frac{1+3i}{5+3i}\times \frac{5-3i}{5-3i}\]

OpenStudy (anonymous):

denomimator is \[25+9=33\]

OpenStudy (anonymous):

numerator is \[5+9+(-3+15)i=14+12i\]

OpenStudy (anonymous):

giving an answer of \[\frac{14+12i}{33}=\frac{14}{33}+\frac{12}{33}i\] in standard form

OpenStudy (radar):

a little short on the denominator.

OpenStudy (anonymous):

oops!

OpenStudy (anonymous):

\[25+9=...\]

OpenStudy (anonymous):

addition never my strong suit !

OpenStudy (radar):

Hi satellite73

OpenStudy (anonymous):

hi radar! thanks for catching my mistake! i guess answer is actually \[\frac{14}{34}+\frac{12}{34}i=\frac{7}{17}+\frac{6}{17}i\] unless i made another mistake